OpenStudy (anonymous):

Verify the identity. cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x

OpenStudy (anonymous):

@keeponbleeding

OpenStudy (anonymous):

@toadsage

OpenStudy (anonymous):

left side: make it up by \[\cos 2x= \cos ^2x-\sin^2x= 1-\sin^2x-\sin^2x= 1 -2\sin^2x\]

OpenStudy (anonymous):

left side too, \[\cos 4x= \cos^2x -\sin^22x= 1-\sin^22x-\sin^22x=1-2\sin^22x\]

OpenStudy (anonymous):

add them together you have left side =right side. done

OpenStudy (anonymous):

add them together how?@Hoa

OpenStudy (anonymous):

cos2x + cos 4x (that is left side) = 1- 2sin^22x+1-2sin^2x = 2 - sin^22x -2sin^2x (right side.) is it not clear?

OpenStudy (anonymous):

ooooo. ok i understand now. thank

OpenStudy (anonymous):